ACTIVITY IN BASIC CALCULUS

BASIC RULES IN DERIVATIVE


 Complete the blanks of the given function below with a number (except 0 and 1) to create your own problem and find the derivative of the function. Show your complete solution to each problem.

 

1. f(x) = -4x5+ ______x-4 - 2468


2. f (x) =____x-3- _____x1/4 - 12x


3.f(x)= ____ "\sqrt[4]{x^3} - \underset{x^6} + \frac{2}{3} x^6"


4.f (x) = "\underset{x^-6}{=} -" ____ x2 + "\sqrt[4]{x}"


1
Expert's answer
2022-03-23T19:14:58-0400

1."f(x)=-4x^5+2x^{-4}-2468"

"f'(x)=-20x^4-8x^{-5}-2468"

2."f(x)=3x^{-3}-4x^{1\/4}-12x"

"f'(x)=-9x^{-4}-x^{-3\/4}-12"

3."f(x)=4\sqrt[4]{x^3} -x^6+\frac{2}{3}x^6"

"f'(x)=3x^{-1\/4}-6x^5+4x^5"

4. "f(x)=x^{-6}-2x^2+\sqrt[4]{x}"

"f'(x)=-6x-4x+\frac{1}{4}x^{-3/4}"